# Relative pose of three calibrated and partially calibrated cameras from four points using virtual correspondences

### v2

@@ 9.4 Timings
In the main paper, we presented run-times of the proposed solvers as well as the state-of-the-art solvers for the relative pose problem of three calibrated cameras (cf. Table 1 (last column) in the main paper). Here we present run-times of solvers for partially calibrated cameras. To measure the run-times of the solvers, we calculated the average run-time of each solver on more than 10k instances of the Sacre Coeur scene of the PhotoTourism dataset [54]. The run-times are reported in Table 3. The experiments were performed on an Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz. The average run-times of the L-based solvers are higher, because we run the network on the CPU and without batching. In general, the implementations of all proposed solvers are not optimized for speed, and we still see room for speeding them up.
9.5 $\delta$-based solvers
We tested our $\delta$-based solvers for different values of $\delta$ and measured their performance. The results of this experiment for the calibrated 4p3v problem on the Sacre Coeur and St. Peters’ Square scenes from the PhotoTourism dataset [54] are reported in Table 4.
In general, there is no common value of the $\delta$ shift that leads to the best results on all datasets. This is expected since the precision of the mean-point correspondence depends on many different factors, e.g., the viewing angles of the cameras, the type of the motion, the depth and spatial distributions of the 3D points, etc. In general, choices of $\delta$ between $0.05$ and $0.15$ resulted in the best results on the tested datasets. For the 4p3v(L±$\pm\delta$), and 4p3v(L±$\pm\delta$init) solvers a bit smaller shifts are preferable, since these solvers are already shifting the mean-point correspondence by a prediction returned by the network. On the tested datasets a good trade-off between the precision and the running time of RANSAC is achieved using $\delta=0.15$ for the 4p3v(M±$\pm\delta$) solver, and $\delta=0.1$ for the 4p3v(L±$\pm\delta$) and 4p3v(L±$\pm\delta$init) solvers.
The results for the partially calibrated 4p3vf problem on the KITTI dataset [17] are reported in Table 5. Similar to the calibrated case, similar performance is observed for various choices of $\delta$. However $\delta=0.1$ seems to be the best choice for all three solvers, namely 4p3vf(M±$\pm\delta$), 4p3vf(L±$\pm\delta$), and 4p3vf(L±$\pm\delta$init).
Designing a strategy for selecting a suitable $\delta$, e.g., based on the shape of the triangle used to compute the virtual correspondence or based on results obtained using different $\delta$-shifts in early iterations of RANSAC, or from a few images from the given sequence of images (if we e.g. have to process a sequence of images that undergo a similar motion) would be an interesting future work.
Sacre Coeur (5000 image triplets) St. Peters’ Square (5000 image triplets)
Estimator AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$ Time (s) $\downarrow$ AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$ Time (s) $\downarrow$
4p3v(M $\pm$ 0.05) 24.35 12.39 44.83 52.43 59.07 13.14 26.28 19.01 20.34 32.95 46.52 12.81
4p3v(M $\pm$ 0.1) 24.42 12.63 44.16 51.85 58.53 13.11 26.16 19.07 19.51 32.44 46.00 13.63
4p3v(M $\pm$ 0.15) 24.49 12.42 44.51 52.27 58.91 13.10 25.96 18.92 19.62 32.53 46.24 12.76
4p3v(M $\pm$ 0.2) 24.24 12.58 44.36 52.15 58.76 13.11 26.32 19.20 19.24 32.15 45.92 12.75
4p3v(M $\pm$ 0.25) 23.80 12.63 44.28 52.12 58.81 13.11 26.63 19.30 19.02 31.90 45.64 12.79
4p3v(L $\pm$ 0.05) 24.37 12.43 44.90 52.49 58.94 13.76 26.00 19.10 20.00 32.83 46.53 14.74
4p3v(L $\pm$ 0.075) 24.32 12.52 44.63 52.36 58.87 15.08 25.86 18.87 20.25 33.01 46.72 14.66
4p3v(L $\pm$ 0.1) 24.36 12.35 44.89 52.48 59.00 13.76 26.12 19.06 19.77 32.52 46.06 13.39
4p3v(L $\pm$ 0.15) 24.09 12.64 44.22 52.24 59.01 13.87 26.20 19.18 19.56 32.31 46.01 13.39
4p3v(L $\pm$ 0.05 init) 24.21 12.42 45.04 52.64 59.30 16.30 26.09 19.07 19.89 32.57 46.16 15.78
4p3v(L $\pm$ 0.075 init) 24.36 12.54 44.44 52.15 58.67 16.36 26.23 18.81 19.97 33.03 46.56 16.06
4p3v(L $\pm$ 0.1 init) 23.95 12.46 44.83 52.78 59.29 15.10 26.09 18.83 19.67 32.68 46.48 14.22
4p3v(L $\pm$ 0.15 init) 24.05 12.48 44.46 52.40 59.01 14.90 25.80 18.96 19.48 32.43 46.00 14.53
Table 4: $\delta$-based solvers: The average and median pose errors in degrees, and Area Under the Recall curve (AUC) thresholded at ${}^{\circ}$, ${}^{\circ}$ and ${}^{\circ}$ on a total of 10k image triplets from two scenes of the PhotoTourism dataset [22] for different values of $\delta$ used in the proposed solvers.
Estimator AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$${}^{\circ}$ Time (s) $\downarrow$
4p3vf(M $\pm$ 0.05) 11.65 10.98 15.73 51.36 72.05 85.31 10.09
4p3vf(M $\pm$ 0.075) 11.61 10.97 16.02 51.59 72.23 85.42 10.09
4p3vf(M $\pm$ 0.1) 11.60 10.97 16.11 51.52 72.12 85.40 10.09
4p3vf(M $\pm$ 0.15) 11.66 10.99 15.74 51.36 72.05 85.27 10.07
4p3vf(M $\pm$ 0.2) 11.63 10.99 15.63 51.30 71.97 85.27 10.11
4p3vf(M $\pm$ 0.25) 11.63 10.99 15.68 51.39 72.12 85.36 10.09
4p3vf(L $\pm$ 0.05) 11.66 11.02 15.38 50.47 71.46 84.98 10.15
4p3vf(L $\pm$ 0.075) 11.66 11.01 15.27 50.53 71.58 85.09 10.15
4p3vf(L $\pm$ 0.1) 11.69 11.01 15.44 50.68 71.55 85.07 10.10
4p3vf(L $\pm$ 0.15) 11.72 11.01 15.31 50.54 71.50 84.95 10.16
4p3vf(L $\pm$ 0.05 init) 11.84 11.05 13.98 49.27 70.45 84.13 10.22
4p3vf(L $\pm$ 0.075 init) 11.77 11.04 14.40 49.61 70.89 84.52 10.20
4p3vf(L $\pm$ 0.1 init) 11.74 11.01 15.30 50.30 71.34 84.84 10.18
4p3vf(L $\pm$ 0.15 init) 11.73 11.03 14.74 50.10 71.23 84.74 10.22
Table 5: $\delta$-based solvers: Results on KITTI dataset [17] for partially calibrated cameras and different values of $\delta$ used in the proposed solvers. The reported statistics are the same as in Tab. 4.
9.6 Essential matrix estimation using a virtual correspondence
AVG (${}^{\circ}$) $\downarrow$ MED (${}^{\circ}$) $\downarrow$ ${}^{\circ}$uparrow ${}^{\circ}$${}^{\circ}$ Time (s) $\downarrow$
5pt 5.04 0.89 63.71 74.45 83.11 0.04
4p(M) 5.53 0.93 61.48 72.19 81.30 0.03
4p(M ±$\pm\delta$) 5.21 0.90 61.80 72.33 81.27 0.02
4p(O) 4.40 0.81 65.30 75.88 84.43 0.03
Table 6: The average and median pose errors in degrees, and Area Under the recall Curve (AUC), thresholded at ${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$, as well as the average run-time (in seconds) on 9,900 image pairs from two scenes of the PhotoTourism dataset [22].
The proposed M-based solvers do not exploit information from all three views when generating a virtual correspondence. They use only information from three point correspondences in two views and are thus not specific to the 3-view scenario considered in the paper. This suggests that the proposed method for generating virtual correspondences can be used to solve other minimal problems using sub-minimal samples inside RANSAC.
Here, we show preliminary results for the well-known 5-point relative pose problem, where we sample only four point correspondences inside RANSAC and generate the $5^{\text{th}}$ correspondence (a) as a mean point of three corresponding points in both images (the 4p(M) solver), (b) by considering two additional points in $\delta$-vicinity of the mean point in the second image, thus generating three different $5^{\text{th}}$ correspondences. The two additional points are selecting using the same strategy as used in the 4p3v(M±$\pm\delta$) solver presented in Sec. 3.1 of the main paper. We call this solver the 4p(M ±$\pm\delta$) solver. We do not consider the L-based solver since our network was trained on data from all three images. However, a similar network can trained only on four points in two views instead of four points in three views.
In our experiments, we tested the solvers on the same two scenes (i.e., Sacre Coeur and St. Peter’s Square) from the PhotoTourism [54] dataset that we used in the main paper. The 4,950 image pairs used for each scene were selected by the Image Matching Challenge 2020 [22]. The challenge also provides point correspondences together with the intrinsic calibration matrices and ground truth camera poses obtained from a COLMAP reconstruction [52].
We run the 5pt solver [38], and the proposed 4p(M) and 4p(M±$\pm\delta$) solvers which, respectively, generate the fifth correspondence by taking the mean of the first three points, or via using the mean point and two points from its $\delta$-vicinity to compensate for noise in the “mean-point” correspondence. We also show the results of the oracle 4p(O) solver with a fifth correspondence that is selected ensuring that it is consistent with the ground truth epipolar geometry. All minimal solvers are used inside GC-RANSAC [3].
The results, over all 9,900 image pairs from both scenes, are reported in Table 6. While the virtual $5^{\text{th}}$ correspondences generated in the proposed 4p(M) and 4p(M±$\pm\delta$) solvers can be noisy, both solvers show only slightly worse results than the 5pt solver, which uses a more accurate $5^{\text{th}}$ correspondence, inside GC-RANSAC. Moreover, even though the running time of the 4p(M±$\pm\delta$) solver is approximately 3×3$\times$3 × slower than the running time of the 5pt and the 4p(M) solvers, the running time of the whole GC-RANSAC is almost 2×2$\times$2 × faster, showing that the 4p(M±$\pm\delta$) solver is able to find a good solution in earlier iterations. This shows the potential for applying such sub-minimal solvers in scenarios with high outlier ratios and noisy data, where sampling one less point may lead to faster run-times and where the higher noise of the $5^{\text{th}}$ correspondence will only have a limited impact.
The results of the oracle 4p(O) solver shows that there is still some space for improvement when predicting the $5^{\text{th}}$ virtual correspondence, e.g., using a learning based method. However, in this case, the situation is more “challenging” than for the 4p3v problem. For the two-view problem, four points define an infinite number of camera poses that can observe these points, i.e., we have an under-constrained problem. This means that while the proposed M-based strategy and the proposed network can return quite good correspondences for some image pairs and scenes, there is no guarantee that it will not result in larger errors for some other image pairs.
10 Future Work
Our work opens up several directions of future research. Here we list some of them:
  • First of all, a detailed analysis on the noise that the mean point correspondences introduce, to provide theoretical results and bounds through uncertainty propagation analysis, would be important to investigate.
  • Designing and training more complex and efficient network architectures for both the 4p3v(L) and 4p3vf(L) solvers can be a path to achieving better performance for the 4p3v and 4p3vf problems. For 4p3vf(L) it would be reasonable to train a separate model because the 6pt relative pose problem has many more degenerate configurations than the 5pt relative pose problem.
  • On the implementation side, one could speed up the $\delta$-based solver by running the multiple 5pt/ 6pt solvers in parallel or generating $\delta$-correspondences only if the mean-point correspondence returns a reasonable number of inliers inside RANSAC. This would significantly reduce the run-times of those solvers. One can speed up the learning-based solvers by running the inference on a GPU instead of a CPU.
  • Designing and testing different sampling techniques inside RANSAC, so that the proposed M-based and L-based solvers would benefit on average from the spatial distribution of points in these samples (e.g., the size and shape of a triangle from which the mean point is computed) might improve both performance and run-time.
  • Applying the proposed method based on virtual correspondences to new camera geometry problems, e.g., the complex rolling shutter relative pose problem for which minimal solvers do not exist but non-minimal solvers using a non-minimal number of points might be possible to implement reasonably efficiently.

### v3

@@ 9.6 Detailed experiments on Real Data
9.7 Oracle solvers
[…]
To obtain upper bounds for the precision that can be achieved by our proposed solvers we consider the following “oracle” solvers for real experiments: The 4p3v(O)/4p3vf(O) solvers use a correct correspondence(s), i.e., a correspondence(s) that satisfies the epipolar constraint for the ground truth relative pose of the first two cameras, as the $5^{\text{th}}$/$6^{\text{th}}$ virtual correspondence between these cameras. Then the 5pt+P3P/6pt+P3P solver is applied to estimate the relative pose of three cameras. The 4p3v(O)/4p3vf(O) solvers thus indicate the maximum precision that our solvers can reach, if they would have been able to predict or infer a precise $5^{\text{th}}$/$6^{\text{th}}$ correspondence from the coordinates of four input correspondences.
Tab. 7 compares the 4p3v(NO) with the 4p3v(O) solver as well as various variants of our M-based and L-based solvers as well as the 4p3v(HC) [19] and 5pt+P3P approaches. Results are shown over all scenes from the PhotoTourism dataset using GC-RANSAC. The 4p3v(NO) solver performs slightly better than 4p3v(O) in terms of pose accuracy. [^6] We observed that when naively applying 4p3v(NO) in RANSAC, RANSAC tends to terminate too early, resulting in reduced pose accuracy. This resulted in the statement in the main paper that the 4p3v(NO) oracle performed worse than our oracles. In order to obtain results comparable with 4p3v(O), we had to adapt RANSAC. While the pose accuracy of 4p3v(NO) provides an upper bound on the performance of the 4p3v(N) solver [35], the run-time observed for 4p3v(NO) is not indicative of the run-time of 4p3v(N). As highlighted above and in the main paper, the 4p3v(N) solver needs to evaluate up to thousands of epipole estimates and is thus significantly slower than its oracle variant (which only evaluates a single epipole). In contrast, our M-based solvers have a run-time comparable to 4p3v(O). In addition, the epipoles used by 4p3v(N) can be rather noisy, depending on how densely the curve is sampled. Even for a rather dense sampling, which increases the run-time of the 4p3v(N) solver, [35] report that their solver often has problems with local minima. The results of the 4p3v(O) solver show that there is room for improvement for our M-based and L-based solvers by developing approaches to generating more accurate virtual correspondences.
9.8 Solver run-times
@@
In this section, we present run-times of the proposed solvers as well as the state-of-the-art solvers for the relative pose problem of three calibrated/partially calibrated cameras. While the main paper reports run-time results for full RANSAC-based estimation, we now report the run-times of the individual solvers outside of RANSAC. To measure the run-times of the solvers, we calculated the average run-time of each solver on more than 10k instances of the Sacre Coeur scene of the PhotoTourism dataset [51]. For calibrated cameras, the run-times are reported in Table 10, and for partially calibrated cameras in Table 11. The experiments were performed on an Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz. The average run-times of the L-based solvers are higher, because we run the network on the CPU and without batching. In general, the implementations of all proposed solvers are not optimized for speed, and we still see room for speeding them up.